During a tornado in 2008 the Peachtree Plaza Westin Hotel in downtown Atlanta suffered damage. Suppose a piece of glass dropped near the top of the hotel falling 215 meters. Ignoring air resistance, what is the initial velocity of the piece of glass?
1
Expert's answer
2020-11-16T07:37:23-0500
"v^2=2gh\\v^2=2(9.8)(215)\\v=65\frac{m}{s}"